OpenStudy (anonymous):

A circle has a diameter of 8 millimeters. What are the circumference and area of the circle? Use three point one four for pi. Round to the nearest tenth. twelve point six millimeters; two hundred one square millimeters twenty-five point one millimeters; two hundred one square millimeters twelve point six millimeters; fifty point two square millimeters twenty-five point one millimeters; fifty point two square millimeters

OpenStudy (anonymous):

This is pretty simple. Circumference=2*pi*r Area=pi*r^2

OpenStudy (anonymous):

Actually Mokeira, it would be A=3.14*4^2

OpenStudy (anonymous):

There's a diameter of 8, but the formula calls for radius squared

OpenStudy (midhun.madhu1987):